A swimming pool can be filled by 5 pumps working for 2 hrs each. If one of the pumps is non-functional, how much extra time would be required to fill the pool?
a) 20% b) 50% c) 25% d) 100%

Answer by ankor@dixie-net.com(22740)

A swimming pool can be filled by 5 pumps working for 2 hrs each.
find the no. of pump-hrs required to fill the pool
5 * 2 = 10 pump-hrs required to fill the pool
:
If one of the pumps is non-functional, how much extra time would be required to fill the pool?
let t = time required with 4 pumps
4t = 10
t = 10/4
t = 2.5 hrs
:
A half hr longer, find the percentage longer
.5/2 * 100 = 25% longer
.